A coil having N turns is wound tightly in the form of a spiral with inner and outer radii a and b respectively.When current I passes through the coil,the magnetic field at centre is??

Grade:12

1 Answers

Mastermind Joker
6 years ago
Consider an element of thickness dr at a distance r from coildn = dr * No of turns per unit length = dr* N/(b-a) Using Biot savart law, dB = μ0Idn/2rdB = μ0INdr/(b-a)r B = integral(μ0INdr/(b-a)r) b/w a and b = μ0INln(b/a)/b-a